OpenStudy (anonymous):

How do i factor 8x^3-27?

OpenStudy (anonymous):

a^3 - b^3 = (a - b) * (a² + ab + b²) a = 2x and b = -3 <=> (2x)^3 - (-3)^3 = (2x - (-3)) * ((2x)² + (2x * (-3)) + (-3)²) = (2x + 3) * (4x² - 6x + 9)

OpenStudy (anonymous):

medal plzz :)

OpenStudy (anonymous):

hopefully u get the process

OpenStudy (anonymous):

Nice @britbrat4290

OpenStudy (goformit100):

You can do (2x + 3) * (4x² - 6x + 9)

OpenStudy (anonymous):

yea u can do it a couple ways

OpenStudy (goformit100):

And that will be your required answer

OpenStudy (anonymous):

ty @Paynesdad :)

OpenStudy (anonymous):

So I can solve (2x + 3) * (4x² - 6x + 9)

OpenStudy (anonymous):

There is no solving it is just factoring it.

OpenStudy (anonymous):

You can't solve something unless it is an equation or an inequality

OpenStudy (anonymous):

How do i come up with the answer? Like its 8x^3-27 but i dont see the GCF so how do i factor now?

OpenStudy (anonymous):

How did i get 2 for the x?

OpenStudy (anonymous):

if u do the math